# Deep source separation of overlapping gravitational-wave signals and non-stationary noise artifacts

### v1

@@ IV.1 Training data and simulation setup
The MBHB waveforms span a mass range of $10^{5}M_{\odot}$ to $10^{6}M_{\odot}$ per binary component, distributed across redshifts from $z=2$ to $z=5$. These sources are modeled using spin-aligned inspiral-merger-ringdown waveforms with higher-order harmonic modes, simulating the final inspiral phase visible to LISA. GBs are generated from a population of compact white-dwarf binaries, covering frequencies between 1 mHz and 10 mHz. In this proof-of-concept study, the individual GB strains are considered to lie within the range of $1\times 10^{-22}$ to $2\times 10^{-21}$. We assume an upper limit of $N_{\text{GB}}=5$ GBs, indicating that each dataset may include up to five time-overlapping GB sources in addition to MBHBs and glitches. However, the actual number could be lower, remains unknown a priori, and must be inferred by the model. Scaling to realistic GB populations is underway and will be the subject of future publications. The instrumental noise follows the expected LISA sensitivity curve, while transient glitches are drawn from a randomized distribution to reflect possible instrumental artifacts.
We train the model with a fixed learning rate of $10^{-3}$, the Adam optimizer, and the loss function of Eq. 5. Training is performed for 25 epochs with a batch size of 16, ensuring convergence across all decoder heads. The training dataset comprises 25,000 different data samples.
@@ IV.3 Model evaluation on unseen test data
Here, $X(f)$ and $\hat{X}(f)$ are the Fourier transforms of $x$ and $\hat{x}$, respectively.
@@ IV.3.4 Resolving individual GBs in the presence of MBHBs and glitches
Finally, we evaluate the performance of the GB decoder in distinguishing individual and overlapping GBs. Figures 9 and 10 compare the true and predicted number of GBs, demonstrating that the model accurately estimates the number of active sources even in the presence of glitches and merging MBHB. This supports the effectiveness of the frequency-bin approach. We notice that the outputs of silent channels are not entirely zero. This is expected, given the decoder design. In future iterations, we plan to refine the GB decoder head by incorporating a gating mechanism that learns to fully suppress inactive channels, minimizing spurious noise when no signal is present in a given bin.
From this perspective, our model is designed with scalability in mind; we train on short data snippets and plan to apply a Demucs-like stitching procedure in future work to process arbitrarily long datasets once the model is trained. This approach involves segmenting long data streams into manageable, overlapping pieces, processing each segment individually, and then combining the results to reconstruct the full signal. This method has been effective demucs and shows promise for application in gravitational wave data analysis. However, it’s important to note that this stitching procedure is not part of the current study and will be explored in a follow-up investigation.
Regarding the estimation of astrophysical source parameters and their posterior distributions, our current approach is centered on source separation rather than direct Bayesian inference. The encoder-decoder framework extracts individual signals from a noisy mixture, producing separated outputs for each source. While this does not inherently yield posterior distributions, the extracted signals can serve as inputs to subsequent inference models. Future work will explore incorporating Bayesian techniques, such as neural density estimators (e.g., normalizing flows), to estimate posterior distributions of source parameters based on the separated signals. This will provide a more comprehensive understanding of uncertainties while leveraging the strengths of our blind source separation framework.
@@ V Conclusion
In this work, we presented a deep learning-based framework for blind source separation in high-dimensional LISA data, addressing overlapping gravitational-wave signals and non-stationary noise artifacts. Inspired by demucs, a model originally designed for audio processing, our approach employs a shared encoder-decoder architecture to disentangle complex signal components directly in a single step, bypassing iterative subtraction techniques, which represents a conceptual shift in methodology. The model isolates individual components dynamically through latent space clustering while remaining scalable to high-density astrophysical populations.
@@
Our current implementation is a proof-of-concept study, that restricts the number of overlapping GBs and excludes complicated EMRI waveforms. Nevertheless, the results demonstrate the potential of deep source separation for LISA data analysis. Notably, even with this simple framework – consisting of a few hundred lines of code and trained on a modest dataset – model inference operates efficiently on a standard laptop within seconds. Building on the presented results, we will explore the framework’s scalability to more complex, large-scale source populations. We are optimistic about its broader applicability.

### v2

@@ IV.1 Training data and simulation setup
MBHBs.
The MBHB waveforms span component masses uniformly sampled in between $10^{5}M_{\odot}$ and $10^{6}M_{\odot}$, with redshifts drawn uniformly in comoving volume over the range $z=2$ to $z=5$. The binaries are assumed to be spin-aligned and non-precessing, and the waveforms include inspiral, merger, and ringdown phases with higher-order harmonics using the IMRPhenomHM approximant.
GBs.
GB signals are modeled as slowly drifting sinusoids to reflect the intrinsic frequency evolution of compact white-dwarf binaries over multi-hour timescales. Their frequencies are drawn uniformly in log-space from 1 mHz to 10 mHz, and the strain amplitudes are sampled uniformly in log-space between $1\times 10^{-22}$ and $2\times 10^{-21}$. For each sample, up to five GBs are included ($N_{\text{GB}}\leq 5$), with the actual number per simulation drawn from a uniform discrete distribution. Sky positions are sampled isotropically, and polarization and inclination angles are drawn uniformly over their natural ranges.
While this setup does not include the full Galactic foreground, which is expected to form a confusion-limited noise floor in LISA’s low-frequency band, it enables us to assess the model’s separation capabilities in interpretable conditions. We acknowledge that a realistic mission scenario will operate near the detection limit for most GBs as spectral resolution improves and that resolving overlapping near-threshold sources will be significantly more challenging. Scaling to such high-density GB populations is underway and will be addressed in future work.
Frequency binning.
The frequency range from 1 mHz to 10 mHz is divided into $K=5$ uniform-width bins, each spanning 1.8 mHz. This bin count matches the maximum number of overlapping GB sources per training sample, providing a clean mapping in which each source can, in principle, be assigned to its own non-overlapping spectral bin. Within each bin, sources are aggregated and treated as a single target during training, and the model learns to reconstruct this frequency-binned representation.
While this design simplifies the problem and enables a proof of concept, it does not reflect the full complexity expected in LISA data. In reality, GB sources will be densely distributed in frequency space, with many overlapping in narrow bands. The current setup thus represents a tractable starting point for demonstrating the feasibility of source separation in moderately crowded conditions. We emphasize that this binning strategy is a first step, and future work will address more realistic scenarios involving higher GB densities and stronger spectral overlap. To support this, we outlined architectural enhancements in the previous section.
Glitches.
Transient glitches are modeled as localized Gaussian bursts, where the amplitude is chosen to yield a broad range of SNRs. Specifically, we target glitches ranging from low-SNR cases that are buried in the quasi-stationary noise to high-SNR transients that can exceed the peak amplitude of the MBHB signal. Each sample includes between 0 and 30 such glitches, with randomly sampled locations.
Noise model.
The instrumental noise follows the standard LISA noise curve, including optical metrology and test mass acceleration noise contributions. It is illustrated in Fig. 3.
Training details.
We train the model using the Adam optimizer with a fixed learning rate of $10^{-3}$, optimizing the loss function defined in Eq. 5. Each model is trained for 25 epochs with a batch size of 16, ensuring convergence across all decoder heads. The training dataset contains 25,000 samples.
Figure 3: LISA noise curves for the TDI A, E, and T channels. We use these profiles to generate colored Gaussian noise in our simulations, ensuring that the synthetic noise matches the expected characteristics of the LISA instrument.
@@ IV.3 Model evaluation on unseen test data
Here, $X(f)$ and $\hat{X}(f)$ are the Fourier transforms of $x$ and $\hat{x}$, respectively. This noise-weighted inner product is standard in gravitational-wave data analysis and reflects the optimal matched filtering statistic under Gaussian noise assumptions. The weighting by $1/S_{n}^{X}(f)$ down weights frequency regions with high noise (low sensitivity) and emphasizes those where the detector is most sensitive. As such, it ensures that waveform agreement is judged in terms of physically relevant distinguishability in the presence of instrumental noise.
@@ IV.3.1 Glitches overlapping with MBHB signal
To further quantify the impact of glitches on MBHB waveform recovery, we simulate datasets that include an MBHB signal, a single instrumental glitch, and stationary LISA-like noise. In each simulation, the MBHB and glitch components are independently normalized to achieve comparable SNRs, ensuring that both contribute similarly to the time-domain mixture. Balancing the SNRs establishes a controlled setting where both the astrophysical signal and the glitch influence the data similarly, preventing trivial cases in which one component dominates. This setup allows us to probe their interference and disentanglement during recovery meaningfully.
We then systematically vary the relative timing between the MBHB coalescence and the glitch, shifting the glitch by an offset ranging from –2 hours to 0 (coalescence time) while keeping the MBHB and noise fixed. At each offset, we generate the noisy TDI mixture, apply the deep source separation model, and compute the noise-weighted match factor between the true and recovered MBHB waveforms. Figure 8 displays the median match factor $M$ as a function of glitch offset (solid red line), along with the 25–75% interquartile range (dark shaded region) and the 5–95% percentile range (light shaded region) over 150,000 randomized glitch–MBHB realizations. The x-axis denotes the glitch offset in hours before coalescence, with the vertical dashed line marking the moment of the MBHB merger.
Figure 8: Impact of glitch timing on MBHB recovery. We simulate mixtures of an MBHB signal, a single glitch, and LISA-like noise, scaling the MBHB and glitch to similar SNRs. In this figure, we set the SNR to 50. The glitch is systematically shifted in time relative to the MBHB coalescence, and the deep source separation model is used to recover the MBHB waveform across all offset configurations. The match factor remains nearly constant, with a slight decline only when the glitch overlaps the merger. Higher SNRs yield higher match values, but the overall behavior remains consistent.
The match factor remains consistently high across most of the inspiral phase, with only a modest decline occurring when the glitch temporally overlaps with the merger. This suggests that the deep source separation model can quite robustly recover MBHB signals. One reason for this robustness lies in the distinct spectral and temporal characteristics of MBHB signals versus glitches. MBHB mergers are coherent chirps, whereas glitches may appear as abrupt, high-frequency bursts or narrow-band transients. The MBHB decoder learns to recognize the typical MBHB waveform morphology in the latent space representations, even in time-overlapping cases.
@@ IV.3.2 Weak MBHB mergers buried in noise
Figure 10 illustrates the relationship between the SNR and the normalized match factor $M$ for MBHB signals recovered from noisy TDI data using the trained deep source separation model. Each gray point represents a single simulation; in total, 150,000 points are shown. The red curves indicate the median match (solid line), the 25–75% interquartile range (dark shaded region), and the 10–95% percentile spread (light shaded region).A vertical line marks the SNR at which the median match first exceeds the accuracy threshold of $M=0.95$, which in this analysis occurs at approximately SNR $\approx 15$. This threshold is chosen for illustrative purposes. A systematic investigation is needed to determine how well waveform parameters can be recovered as a function of the match factor. We present preliminary results on this question at the end of this section, though we note that the primary focus of this paper is on deep source separation rather than source parameter inference.
Figure 10: Match factor $M$ as a function of SNR for quiet MBHB signals recovered from noisy TDI data. Gray points show individual simulations, and red curves indicate the median and percentile bands. The vertical line marks where the median first exceeds $M=0.95$. The value is arbitrarily chosen for visualization. SNRs are computed empirically, treating glitches and GBs as noise.
To probe match performance across a broad SNR range, we generate multiple amplitude-scaled MBHB signals per realization, resulting in effective SNRs spanning from 5 to 100. In this setup, GBs and transient glitches are treated as part of the effective noise background, as they hinder the accurate recovery of the MBHB waveform. The empirical SNR for each signal is calculated by estimating the PSD from the effective noise background using Welch’s method. The inset zooms into the high-SNR regime. The tightening of percentile bands with increasing SNR indicates improved reconstruction stability, while the saturation of the median match near unity confirms the model’s effectiveness in extracting MBHB signals under favorable conditions.
@@ IV.3.3 Quiet periods with no MBHB mergers
To further evaluate the reliability of the MBHB decoder in distinguishing true astrophysical signals from spurious activations, we analyze the empirical cumulative distribution functions (ECDFs) of the decoder output power. Specifically, we consider the squared $\ell_{2}$-norm of the decoder’s output in the MBHB channel,
$P=\sum_{t}\hat{x}^{2}_{\mathrm{MBHB}}(t),$ (13)
where $\hat{x}_{\mathrm{MBHB}}(t)$ is the predicted strain at time $t$. The ECDF, defined as
$\mathrm{ECDF}(P_{0})=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{P_{i}\leq P_{0}\}},$ (14)
provides the cumulative fraction of samples whose decoder output power does not exceed a given threshold $P_{0}$. Intuitively, the ECDF tells us for any power value, what fraction of decoder outputs were smaller than or equal to that value.
We compare ECDFs for two types of simulated LISA data: one containing only instrument noise and glitches (”No MBHB”), and one in which a MBHB signal is injected (”With MBHB”). The gray curve in Fig. 11 shows the ECDF of decoder power in the absence of a true signal. As expected, the curve rises steeply and saturates at very low power levels, indicating that the decoder remains largely inactive when no MBHB is present. The red curve shows the ECDF for the same decoder when an MBHB signal is included, resulting in a markedly slower rise toward unity and significantly higher output power, reflecting strong decoder activation in response to astrophysical signals. In this case, we again use 150,000 simulation samples, with MBHB parameters randomly varied across the simulations.
Figure 11: ECDFs of the MBHB decoder output power, defined as the squared $\ell_{2}$-norm of the predicted strain. The gray curve shows results for glitch+noise-only inputs, while the red curve includes MBHB signals with varying parameters across simulations. The steep rise of the gray curve indicates minimal decoder activity during quiet periods. The small hallucination area and large ECDF mass confirm the decoder’s low false-positive rate.
To quantify potential spurious responses, we define the hallucination area as the area between the actual ECDF for glitch+noise and the ideal ECDF that would jump directly to one at $P=0$:
[…]
This area captures the total ”excess activation” of the decoder during quiet periods. A small hallucination area implies that the decoder rarely outputs significant power when no MBHB is present, indicating low false-positive risk. In our test data, we find $A_{\mathrm{hall}}=6.88\times 10^{-5}$.
In contrast, we define the ECDF mass as the area under the ECDF curve itself:
[…]
which reflects how quickly the decoder output accumulates across the dataset. A high ECDF mass corresponds to decoder outputs clustering near zero – the ideal behavior when no MBHB is present. For the glitch+noise case shown, the ECDF mass is approximately $2.83\times 10^{2}$.
Together, the separation between the red and gray curves, the small hallucination area, and the high ECDF mass all indicate that the MBHB decoder is well-calibrated and reliably silent during quiet periods while remaining sensitive to true astrophysical signals.
@@ IV.3.4 Resolving individual GBs in the presence of MBHBs and glitches
Finally, we evaluate the performance of the GB decoder in distinguishing individual and overlapping GBs. Figures 12 and 13 compare the true and predicted number of GBs, demonstrating that the model accurately estimates the number of active sources even in the presence of glitches and merging MBHB. This supports the effectiveness of the frequency-bin approach. We notice that the outputs of silent channels are not entirely zero. This is expected, given the decoder design. In future iterations, we plan to refine the GB decoder head by incorporating a gating mechanism that learns to fully suppress inactive channels, minimizing spurious noise when no signal is present in a given bin.
@@
To quantitatively assess the separation performance for GB signals, we perform an ensemble study analogous to Fig. 10 using 150,000 synthetic LISA simulations. Each simulation includes a single variable-strength GB signal injected into a fixed realization of LISA-like noise, possibly including instrumental glitches and MBHBs. The GB signal is scaled using a range of amplitude factors, resulting in target signal-to-noise ratios (SNRs) between 1 and 100. For each mixture, the deep source separation model is applied to recover the GB waveform. Then, for each recovered GB waveform, we compute the match factor between the predicted and true signal as a function of the SNR. The results are given in Fig. Fig. 14. The figure shows the distribution of match values across the ensemble, including the median (red line), interquartile range (dark shaded region), and the 5–95% percentile range (light shaded region). The background density plot visualizes the individual scatter of simulations. At low SNRs, performance is limited, as expected, but the median match improves rapidly and exceeds 0.95 once the SNR reaches approximately 10. In contrast to the analogous analysis performed with the MBHB decoder, we observe a decline in the median match factor for GB signals at high SNRs. This decline reflects the model’s behavior when confronted with out-of-distribution signals rather than an actual performance limitation.
Note that a partial reason for the lower reconstruction accuracy of GBs compared to MBHBs and glitches lies in the relative amplitude differences between source types. Since the total loss is computed as a sum of equally weighted MSE terms, higher-amplitude sources like MBHBs tend to dominate the optimization. As a result, lower-amplitude GBs contribute less to the gradient signal and may be underfit. To address this imbalance, future work will explore adaptive loss weighting schemes, where a secondary network dynamically estimates task-specific weights based on source uncertainty or signal characteristics, allowing the model to balance reconstruction fidelity more uniformly across all components.
Figure 14: Match factor $M$ as a function of SNR for isolated GB signals recovered from noisy TDI data. Gray points represent individual simulations, while red curves show the median and percentile ranges. The vertical line marks the SNR at which the median match first exceeds $M=0.95$, a threshold chosen for visualization purposes. The drop in performance at high SNR reflects the model’s response to out-of-distribution inputs.
From this perspective, our model is designed with scalability in mind; we train on short data snippets and plan to apply a Demucs-like stitching procedure in future work to process arbitrarily long datasets once the model is trained. This approach involves segmenting long data streams into manageable, overlapping pieces, processing each segment individually, and then combining the results to reconstruct the full signal. This method has been effective in demucs and shows promise for application in gravitational wave data analysis. However, it’s important to note that this stitching procedure is not part of the current study and will be explored in a follow-up investigation.
IV.4 Non-stationarity and data gaps in long-duration inference
In our current setup, the model is trained and evaluated on short data snippets of 2 hours in duration. Over such timescales, the Galactic binary population can be approximated as quasi-stationary since the individual sources evolve slowly and the overall structure of the foreground remains largely unchanged. Consequently, the training data – and the latent representations learned by the encoder – reflect only the stationary characteristics of the foreground within each segment. During inference, however, the model will eventually be applied segment-wise to arbitrarily long time series, and the outputs are combined using the aforementioned stitching procedure. This allows slowly varying foreground effects – such as Doppler modulation and frequency evolution – to emerge naturally in the reconstructed outputs.
We also note that the proposed framework is compatible with data gaps. Since each segment is processed independently, any valid data before and after a gap can be separately analyzed and reconstructed, with the gap manifesting as a discontinuity in the stitched output. If higher continuity is desired, established LISA gap mitigation strategies could be integrated at the preprocessing or postprocessing level, including time-domain gating, frequency-domain likelihood adaptation, or statistical inpainting methods. Alternatively, exposing the model to artificially gapped training data may enhance robustness in the presence of real interruptions. These directions offer promising extensions to improve the model’s applicability to realistic mission scenarios.
IV.5 Bayesian inference from separated signals
A natural extension of the present framework involves estimating astrophysical source parameters from the separated signals. While the current model is designed for blind source separation, the resulting waveforms – or alternatively, their latent representations – could serve as inputs to Bayesian inference techniques aimed at recovering posterior distributions over source parameters. Simulation-based inference (SBI) offers a particularly promising approach in this context. By training neural density estimators on synthetic populations, one can learn a mapping from reconstructed signals to posterior distributions. When using the outputs of the decoder as inputs to the inference model, the SBI framework becomes implicitly aware of the separator’s characteristics as it learns to account for any distortions introduced by the encoder-decoder architecture. This separator-aware formulation would ideally allow the posterior estimator to remain well-calibrated even when the signal reconstruction is imperfect.
An alternative or complementary direction involves applying classical MCMC sampling directly to the raw TDI data. This remains the most principled approach when a well-defined likelihood function is available, especially since the noise characteristics are reliably known only for the original TDI data – not for the decoded outputs. While such inference is computationally intensive, it could benefit from incorporating signal estimates from the separator as a preprocessing step: parameter point estimates obtained from the decoder outputs – via matched filtering optimization or regression – could be used to narrow the prior range, thereby accelerating burn-in and improving convergence behavior.
Looking ahead, the modular structure of the present framework naturally supports such extensions. In the present work, we take a first step in this direction by applying SBI to MBHB signals reconstructed from the deep source separation framework. We focus on two representative cases: the high-match MBHB example of Fig. 6, and the lower-match MBHB reconstruction of Fig. 7(c). For now, the inference task is limited to sky localization.
We follow a two-step strategy. In the first step, we train an SBI model on true (injection-level) MBHB waveforms, enabling the neural density estimator to learn posteriors in an idealized, distortion-free setting. This model is then used to infer posteriors for both the true injected waveforms and the corresponding recovered waveforms from the separator. Comparing the two outputs reveals how separation accuracy affects inference: we expect that the high-match example yields a nearly identical posterior to the true case, while the lower-match example shows deviations due to waveform distortion.
In the second step, we train a new SBI model with equal architecture using the recovered waveforms themselves as input. This model becomes separator-aware, learning to map distorted inputs to calibrated posteriors despite imperfections in signal reconstruction. We then compare its output – obtained using the recovered waveform – to the output of the original SBI model applied to the true injection. In the ideal case, both approaches yield comparable posteriors, demonstrating that separator-aware inference can recover the correct parameter distribution even in the presence of encoding and decoding artifacts.
A schematic overview of this pipeline is provided in Fig. 16. While we focus here on MBHB sky localization, the methodology generalizes to other parameters and source classes.
[…]
Figure 16: End-to-end separator-inference pipeline. Noisy LISA data is processed by the deep source separation model that disentangles the input into MBHB, Galactic binary, and glitch components. Each recovered signal is then passed to a source-specific simulation-based inference model to estimate physical parameters and infer posterior distributions.
The SBI model is built using the sequential neural posterior estimation framework provided by the sbi library [85]. Posterior distributions are modeled using neural spline flows, which offer a flexible and expressive class of density estimators. The prior is chosen as a uniform box distribution over the sky parameters $\lambda$ and $\beta$. Importantly, the training of the SBI model is fully decoupled from the training of the deep source separation network. We use 50,000 simulated examples to train the SBI model. This setup is intentionally kept simple and is not tuned for performance optimization; the goal is to facilitate controlled comparisons and assess the impact of reconstruction quality on downstream parameter estimation.
Figure 17 illustrates the result of the first-stage inference for the high-match case of Fig. 6. Here, we apply the SBI model trained on true injections to both the true MBHB waveform and the corresponding waveform recovered by the separator. The two posterior distributions are nearly identical in both sky latitude and longitude, demonstrating that when the reconstructed signal closely matches the true waveform, the downstream inference remains virtually unaffected.
In contrast, Fig. 18 shows the same evaluation procedure applied to the lower-match example of Fig. 7(c), where the MBHB waveform is partially distorted. In this case, the posteriors obtained from the true and recovered waveforms begin to diverge, particularly in sky longitude. This indicates that the inference quality degrades gracefully under moderate distortion and suggests room for improvement in modeling separator-induced uncertainty.
To mitigate this, we train the SBI model on the recovered waveforms instead of true injections. As shown in Fig. 19, this separator-aware SBI framework successfully compensates for distortions introduced during source separation. The posterior obtained using the recovered waveform now aligns closely with the one from the true injection evaluated by the original SBI model. This demonstrates that the inference network can adapt to the reconstruction artifacts it is exposed to during training, effectively learning to ’undo’ the distortions introduced by the encoder-decoder model.
This initial integration of deep source separation with SBI serves as a proof of concept and outlook. Future work will extend this approach to a broader population of sources, incorporating full parameter estimation and systematic robustness studies across the LISA sensitivity range.
Figure 17: Posterior distributions from the SBI model trained on true injections, evaluated on both the true waveform and the high-match reconstructed waveform of Fig. 6. The two posteriors are nearly identical, indicating that accurate reconstruction preserves inference quality.
Figure 18: Posteriors obtained using the SBI model trained on true injections for the lower-match case of Fig. 7(c). The distributions begin to diverge, particularly in sky longitude, reflecting the impact of signal distortion on downstream inference.
Figure 19: Posterior from the separator-aware SBI model trained directly on recovered waveforms, corresponding to the lower-match example shown in Fig. 7(c). Despite distortions in the input, the inferred posterior matches the true-injection result from the standard SBI, demonstrating that the re-trained network has learned to compensate for reconstruction artifacts.
@@ V Conclusion
We presented a deep learning-based framework for blind source separation in high-dimensional LISA data, addressing overlapping gravitational-wave signals and non-stationary noise artifacts. Inspired by demucs, a model originally designed for audio processing, our approach employs a shared encoder-decoder architecture to disentangle complex signal components directly in a single step, bypassing iterative subtraction techniques, which represents a conceptual shift in methodology. The model isolates individual components dynamically through latent space clustering while remaining scalable to high-density astrophysical populations. A remaining limitation is that the current study restricts the degree of frequency overlap among individual GB sources, which will be addressed in future work.
@@
Our current implementation is a proof-of-concept study, that restricts the number of overlapping GBs and excludes complicated EMRI waveforms. Nevertheless, the results demonstrate the potential of deep source separation for LISA data analysis. Notably, even with this simple framework – consisting of a few hundred lines of code and trained on a modest dataset – model inference operates efficiently on a standard laptop within seconds. Further training and implementation details are provided in Appendix A. Building on the presented results, we will explore the framework’s scalability to more complex, large-scale source populations. We are optimistic about its broader applicability.
In parallel, we have taken initial steps to integrate SBI into the analysis pipeline. Leveraging neural density estimators trained on synthetic waveforms, SBI enables direct posterior estimation from the separated signals, even in the presence of reconstruction artifacts. This approach is particularly valuable for enabling fast, calibrated parameter estimation without requiring a full likelihood model. As shown in our proof-of-concept results, separator-aware inference models can learn to compensate for distortions introduced by the encoder-decoder architecture, maintaining robust performance across a range of reconstruction qualities. Looking ahead, we envision this integration as a foundation for end-to-end gravitational-wave analysis pipelines, where source separation and parameter inference are tightly coupled within a unified learning-based framework.
@@
Appendix A Training setup and model complexity
The total number of trainable parameters in the Demucs-style multi-source separation model is approximately 2.79 million. This figure reflects the combined weights of the shared encoder, bottleneck, and three decoders targeting MBHBs, glitches, and Galactic binaries (GBs), respectively. The GB decoder outputs one signal per predefined frequency bin; in this work, we use five such bins.
The deep source separation model was trained on a dataset of 25,000 simulated time-domain mixtures, each comprising MBHB signals, glitches, and multiple GB sources embedded in instrumental noise. Each training segment corresponds to a 2-hour duration.
Training was conducted on a MacBook Pro with an M2 Max chip and 32 GB of unified memory, using PyTorch. The model was trained for 25 epochs, requiring approximately 4 hours.
This configuration achieved satisfactory source separation performance for initial evaluations. Larger-scale experiments using GPU-accelerated hardware are underway to assess scalability with increased training data and more expressive model architectures.